The Advanced Photon Source
a U.S. Department of Energy Office of Science User Facility

APS-UES Meeting Minutes - 4/15/2015

Attendees: Bruce Glagola, Nena Moonier, Bruno Fieramosca, Wendy VanWingeren, Shane Flood, and Patti Pedergnana

Bruce Glagola

  • The shutdown starts on April 28th.
  • Today is Tech Tuesday, Bruno will go to 431, Patti to 433, and Shane to 438.

Shane Flood

  • The paging system now goes to the group cell phone.The phone is more reliable.As of now there are four notifications coming through with each page, two on the pager and two on the phone.
  • The FC contact card will have our contact phone number and online request page link.

Patti Pedergnana

  • The XSD-SSM group Three-Year Safety Review is tomorrow.
  • Tecknit is taking the 35-ID-C door 2 off of the station starting today.
  • The 434 residents are aware of the LTHW work at 15-ID during the shutdown.There have been no concerns or comments from the residents.

Wendy VanWingeren

  • Stubbs is starting rad samples in 13-BM-C tomorrow, but setting up today.
  • Don Brown will be running rad samples in 1-ID-C tomorrow.
  • Nenad is running rad samples in 16-BM-B.
  • Kemner is starting rad at 10-BM tomorrow and at 10-ID on Friday.

Nena Moonier

  • Sector 13 is done running their brain samples.
  • Sector 23’s brain samples are secured in the soil cabinet.The users are moving these samples onto Sector 2 next week.

Bruno Fieramosca

  • Tecknit will remove the outboard wall at 8-BM-B.The project was moved up to an April 20th start date.
  • MOM-VAC is installing a new pinhole camera at 35-ID-C.The FEV is closed and Bruno has removed APS enable from the beamline per SI request.Bruno will have Clay follow-up on this work during the evening shift.